Operations Report – November 14th

SOL: 2
Name of person filing report: Donald "Duckie" Jacques
Non-nominal systems: None
Notes on non-nominal systems:
Spirit rover used: NO
Hours: (before EVA)
Beginning charge: (Before EVA)
Ending charge: (On return from EVA, before recharging)
Currently charging:
Opportunity rover used: NO
Beginning charge:
Ending charge:
Currently charging:
Curiosity rover used: YES
Hours: 2780
Beginning charge: 100%
Ending charge: 35%
Currently charging: YES
Perseverance rover used: YES
Hours: 2077
Beginning charge: 100 %
Ending charge: 36%
Currently charging: YES
General notes on rovers:
Summary of Hab operations: put notes here
WATER USE: 114.1 gallons
Water (static tank): 438.9 gallons
Static tank pipe heater (on or off): ON
Static tank heater (On or off) ON
Toilet tank emptied: Yes
Summary of internet: put notes here
Summary of suits and radios: During DOFF of sim suit #1 the retaining ring holding the orange collar came off. Was able to restore the ring, and will inspect others tonight to assure all are secure.
Summary of GreenHab operations: put notes here
WATER USE: gallons
Heater: On
Supplemental light: Off
Harvest: None
Summary of Science Dome operations: Guilliome continued his work with his purple bacteria
Dual split: Off
Summary of RAM operations: (Tools used, work done)
Summary of any observatory issues: Not used
Summary of health and safety issues: None Noted
Questions, concerns and requests to Mission Support: put notes here

GreenHab Report – November 14th

Crew 286 GreenHab Report 14-11-2023
Sol: 2
GreenHab Officer: Donald "Duckie" Jacques
Environmental control: heater On, door closed.
Average temperatures: 68 deg F at 9 am, 109 deg F at mid day, 72 deg F at 5pm
Hours of supplemental light: None
Daily water usage for crops: 10 gallons
Daily water usage for research and/or other purposes: 4 gallons transferred to the science dome for use in the sink.
Water in Blue Tank (200 gallon capacity): 186 gallons
Time(s) of watering for crops: 11 am
Changes to crops:planted additional items in the large green tub, and in the two long trays.
Narrative: I noted that even though the moisture meter registered that every pot was "Wet" I found all very dry, and crumbly to the touch. Per instructions, I prepared fertilizer water for all plants, planted additional items, and a final light watering for newly placed seeds. I was able to complete 1/2 of my targeted planting today, and plan to complete the other half tomorrow.
Harvest: (include which crop and mass in grams) None

EVA Report – November 14th

Crew 286 EVA #3 report 14-NOV-2023
EVA # 3
Author: Roger Gilbertson
Purpose of EVA: To explore and test sites for future hi-res LIDAR scanning and outdoor music performance.
Start time for EVA: 2:30 pm
End time for EVA: 4:35 pm
Narrative: Getting prepared was more efficient, thanks to lessons leaned from EVA 1 and 2, and the newly drafted EVA checklist. GPS monitors were worn by two crew members. The crew reported what seemed to be a rapid drop in remaining vehicle charge during their drive south. Curiosity was at 65% and Perseverance at 62% even with modest driving speeds. Due to that, they stopped their drive in the vicinity of Zubrin’s Head, and short of the intended Barrainca Butte walking path.
The afternoon EVA 3 departed about 30 minutes late and unfortunately the crew “assumed" a 30 minute extension without specifically requesting one from Mission Support. The EVA drove south on Cow Dung Road, noting a rapid drop in both vehicle’s state of charge. Communications became intermittent due to increasing distance from the repeater. One EVA crew member’s radio seemed to stop working, so they resorted to remaining close together in order to use gestures and shouting. CapCom was not notified of the radio problem until after the end of EVA.
Photo were captured using iPhone and documentary video was captured by professional Panasonic EVA-1, with audio recording via separate a Zoom H8 microphone on a stand, which also recorded sound of the radio coms. LIDAR scans were performed near the vicinity of Zubrin’s Head. A series of scans were collected using a iPhone 13 and NeRF (Neural Radiance Field) software to build 3D models.
The crew returned to the habitat at precisely 4:30 and entered the airlock at 4:35.

Destination: Barrainca Butte (not achieved), Zubrin’s Head reached.
Coordinates (use UTM WSG 84): Barrainca Butte (518799,4250996)
EVA Participants:
• Scott Beibin
• Liz Cole
• Hugo Saugier

Road(s) and routes per MDRS Map: South on Cow Dung Road, to area of Zubrin’s Head (4247500, 520000), park and perform scans in the area. Retrace route back to Hab.
Mode of travel: Driving and walking

Vehicles you will be using: Curiosity, Perseverance

Copyright © The Mars Society. All rights reserved. | Main Site